OpenStudy (anonymous):

@whpalmer4 will you please remind me how to convert mm to inches via cm?

OpenStudy (whpalmer4):

\[10\text{ mm} = 1\text{ cm}\]Divide both sides by 10 mm to get \[\frac{10\text { mm}}{10\text{ mm}} = \frac{1\text{ cm}}{10\text{ mm}}\]\[1 = \frac{1\text{ cm}}{10\text{ mm}}\] \[1 \text{ inch} = 2.54\text{ cm}\]Divide both sides by 2.54 cm to get \[\frac{1\text{ inch}}{2.54\text{ cm}} = 1\] Now we can always multiply by 1 without changing the value, so we have two conversion fractions that we can multiply or divide as needed to convert between inches, cm, and mm. \[35\text{ mm} * \frac{1\text{ cm}}{10\text{ mm}}*\frac{1\text{ inch}}{2.54\text{ cm}} = 35\cancel{\text{ mm}} * \frac{1\cancel{\text{ cm}}}{10\cancel{\text{ mm}}}*\frac{1\text{ inch}}{2.54\cancel{\text{ cm}}}\]\[\qquad = \frac{35*1}{10*2.54}\text{ in} \approx 1.38\text{ in}\]

OpenStudy (anonymous):

so basically we can slide the decimal forward one place in the cm number then divide mm by cm, right?

OpenStudy (anonymous):

if 2.54 cm ≈ 1in, then 25.4 mm also ≈ 1in?

OpenStudy (whpalmer4):

well, to convert between metric units, yes, you just move the decimal point around. But it's much safer to do it in this fashion because it won't let you divide when you should have multiplied or shift the decimal point left when you should have shifted it right, etc. Yes. 1 in = 2.54 cm = 25.4 mm = 0.0254 m = ...

OpenStudy (whpalmer4):

2.54 cm = 1 inch exactly. I put in the approximation because the result I got was actually 1.37795, not 1.38.

OpenStudy (anonymous):

Why can't I have *you* as my professor instead of this inept half-arsed digital prof I have now?

OpenStudy (whpalmer4):

I'm reminded of a comment that the late conductor Leonard Bernstein made to his friend and fellow composer Ned Rorem: "The trouble with you and me, Ned, is that we want everyone in the world to personally love us, and of course that's impossible; you just don't meet everyone in the world." :-)

OpenStudy (anonymous):

Turns out, my son has the same prof but he takes her class face-to-face. He says she is lazy, doesn't know what she's doing and doesn't seem to give a fig for the students. I am taking her class as a DL. She's pitiful and I wish this were not my final semester of this degree.

OpenStudy (anonymous):

I like your personality, but I truly appreciate your explanations. *You* make it make sense.

OpenStudy (whpalmer4):

Personally, I think it would be very challenging to be a teacher who had to teach "all comers" regardless of ability, attitude, etc.

OpenStudy (whpalmer4):

If you get to pick and choose your students, it's a lot more fun.

OpenStudy (anonymous):

Darlin, if you were to be taking this class as a DL, you'd come to the same conclusion my son and I have. You'd most definitely notice that she is not instructor material. She doesn't teach. She throws a list of problems at you and says "go to it. and btw, go do it on mymathlab where one may or may not be able to use the digital text!"

OpenStudy (anonymous):

I want to *LEARN* this, not just regurgitate a correct answer and forget it 10 minutes after the assignment is complete

OpenStudy (whpalmer4):

People like you are why I enjoy helping out here. There's a lot of cool stuff to learn, and it doesn't have to be drudgery.

OpenStudy (whpalmer4):

And it's a fun challenge to try to get people to realize that "hey, I CAN do this stuff" even though they've always thought "I suck at math" or "I just can't understand this"...

OpenStudy (anonymous):

despite this instructor, i still love math. i wish i could keep learning higher and higher math. so much fun, so little cash!

OpenStudy (whpalmer4):

It's a bit of a puzzle of sorts — can I figure out where the instruction they've gotten so far has been defective and fix it? You can keep learning, and it doesn't have to cost you anything.

OpenStudy (anonymous):

it has taken me 20 years to finally understand why all degrees require a math course: to teach us critical thinking skills.

OpenStudy (anonymous):

so how do i teach myself calculus, trig, quantum physics?

OpenStudy (anonymous):

etc

OpenStudy (whpalmer4):

I'm fond of saying that the most important lessons from junior/senior high school math are not the Pythagorean theorem, or the formula for the solution of a quadratic, or how many sides are on a pentagon — it's the ability to read and work carefully, with attention to all the little details.

OpenStudy (anonymous):

GRRRRRRRRRR! This hooey blinking IDIOT teacher doesn't know how to convert kg to g. The instructor-created question states: Convert 1946 kg to grams. Correct answer is 1,946,000 because you multiply kg by 1000 to get g. Her given answer? 1.946 g

OpenStudy (whpalmer4):

another site that is very good for math is the Khan Academy website. Calculus: http://freevideolectures.com/University/Khan-Academy Trigonometry: https://www.khanacademy.org/math/trigonometry

OpenStudy (whpalmer4):

Another possibility is that the intended problem was convert 1946 g to kg, then the numeric portion of the answer is right...

OpenStudy (whpalmer4):

I hate to chat and run, but I must go fetch some children. The wicked witch and I are having a dinner party ;-)

OpenStudy (anonymous):

The problem is, I am having to answer what is written, not what she intends. I am half tempted to send an email to her department chair. This is heartbreaking. My degree and future is in this woman's hands and she's giving incorrect info.

OpenStudy (anonymous):

Oh! Bon appetit and thankee-sai for being here when I needed you most

OpenStudy (anonymous):

BTW, children taste best as shish-kebabs. j/s...

OpenStudy (whpalmer4):

(I was kidding about the dinner party, and none of the kids are named Hansel or Gretel :-) Just write out your answers fully to make it clear you know what you are doing is correct. Makes it easier if you need to appeal the "wrong" answers if you can conclusively demonstrate that they aren't wrong.

OpenStudy (anonymous):

In this case, the teacher IS wrong!

OpenStudy (anonymous):

I've been doing metric conversions for over 30 years now.

OpenStudy (anonymous):

Yes, I know I asked about cm to mm....brain fart :P

OpenStudy (whpalmer4):

Oh, I agree — just make sure you dot the i's and cross the t's so that you don't leave any room for saying "well, your answer isn't correct either"

OpenStudy (anonymous):

I respectfully and politely show my work when I confront her. I do my best to be as professional as possible just in case there is a dispute and the communication is read by someone other than myself and her

OpenStudy (anonymous):

Do you know your mathematical abbreviations?

OpenStudy (anonymous):

Teacher asked: Many quantities have customary symbols. What is the U.S. customary symbol for length? a) in b) m c) ft d) s

OpenStudy (anonymous):

which one would *you* choose?

OpenStudy (anonymous):

I know that in = inches, m = meters, ft = feet, but what does s =?

OpenStudy (anonymous):

And her given correct answer is s.

OpenStudy (whpalmer4):

in my experience s most often abbreviates seconds, as in m/s, ft/s, etc. I am unable to think of a common unit of length whose abbreviation is s. I'm sure you'll appreciate this :-) http://www.wolframalpha.com/input/?i=convert+s+to+in